OpenStudy (anonymous):

x= -b +or- sq.root b sq. - 4ac all divided by 2a (sorry if this doesnt make sense)

OpenStudy (anonymous):

thats quadratic formula

OpenStudy (anonymous):

how do you solve it??

OpenStudy (anonymous):

any specific question ?

OpenStudy (anonymous):

Compare your quadratic equation with \(ax^2+bx+c=0\) find a,b,c then the two roots of x are: \(\huge{x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

OpenStudy (anonymous):

oh ok i get it now, thanks

OpenStudy (anonymous):

welcome ^_^